40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 10 03:05:09 2014, ericq, Update, LSC, Which side of optical spring are we on? comparison.pdf
    Reply  Fri Oct 10 03:25:11 2014, Jenne, Update, LSC, Which side of optical spring are we on (simulation) 10x
    Reply  Mon Oct 13 12:01:28 2014, ericq, Update, LSC, Which side of optical spring are we on? AOInjection_SqrtInv_REFLDC.pngAOinjection_SqrtInv_REFLDC.xml.zip
       Reply  Mon Oct 13 21:20:56 2014, Jenne, Update, LSC, Which side of optical spring are we on? (No progress) 
          Reply  Mon Oct 13 21:59:47 2014, rana, Update, Computer Scripts / Programs, Which side of optical spring are we on? (No progress) 
             Reply  Tue Oct 14 17:38:31 2014, ericq, Update, Computer Scripts / Programs, Rossa and Allegra wiped, Ubuntu 12.04 installed 
                Reply  Tue Oct 14 23:44:42 2014, diego, Update, Computer Scripts / Programs, Rossa and Allegra wiped, Ubuntu 12.04 installed 
                   Reply  Wed Oct 15 17:18:10 2014, Jenne, Update, Computer Scripts / Programs, Dataviewer fix with Ubuntu 12.04 
          Reply  Wed Oct 15 02:58:03 2014, Jenne, Update, LSC, Which side of optical spring are we on?  PRFPMI_MC2injection_PosCARMoffset_REFLDC_TRX_response_14Oct2014_TRX.pdfPRFPMI_MC2injection_PosCARMoffset_REFLDC_TRX_response_14Oct2014_REFLDC.pdfPRFPMI_MC2injection_NegCARMoffset_REFLDC_TRX_response_14Oct2014_TRX.pdfPRFPMI_MC2injection_NegCARMoffset_REFLDC_TRX_response_14Oct2014_REFLDC.pdf
             Reply  Wed Oct 15 19:56:38 2014, Jenne, Update, LSC, Which side of optical spring are we on? Meas vs Model AntispringModel_NegOffsetMeas_comparison.pdfSpringModel_NegOffsetMeas_comparison.pdf
Message ID: 10603     Entry time: Mon Oct 13 21:20:56 2014     In reply to: 10598     Reply to this: 10604   10607
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Which side of optical spring are we on? (No progress) 

[Jenne, Diego]

In order to distinguish between the spring and antispring sides of the CARM resonance, we need to have transfer function measurements down to at least 100 Hz (although lower is better). 

We tried to get some transfer functions the same way Q did, but noticed that (a) we couldn't get any low frequency coherence, and (b) that when we increased the amplitude of the white (well, lowpass at 5kHz) noise, the coherence between the AO injection and REFL DC went down.  Not clear why this is.

Anyhow, we tried taking good ol' fashioned swept sine transfer functions, although eventually the lightbulb came on that the AO path has a highpass in it.  Duh, Jenne.  So, we started trying to actuate on MC2 position rather than the AO path laser frequency.  We didn't get too far though before El Salvadore decided to have a few 7.4 earthquakes.  We're bored of aftershocks knocking us out of lock, so we're going to come back to this tomorrow.

 

ELOG V3.1.3-